# Better RAG using Relevant Information Gain

### v1

@@ Abstract
Better RAG using Relevant Information Gain
Marc Pickett                        Jeremy Hartman
Ayan Kumar Bhowmick                        Raquib-ul Alam                        Aditya Vempaty
Emergence AI Contact: mpickett@emergence.ai
@@ 1 Introduction
A limitation of transformer-based Large Language Models (LLMs) is that the number of tokens is bounded by the transformer’s context window, which is typically in the thousands. This is often insufficient for representing large texts, such as novels and corporate documentation. A common way to mitigate this constraint is via retrieval augmented generation (RAG), in which a relatively small subset of relevant passages are retrieved from a larger database and inserted into an LLM’s context window Gao et al. (2024). Typically, this process involves applying a similarity metric, such as cosine similarity, to (precomputed) embeddings of passages and the embedding of a query. Using this metric, many systems then use K-nearest-neighbors or a fast approximation with a vector database such as FAISS Douze et al. (2024). Importantly, K-nearest-neighbors Bijalwan et al. (2014) and related methods (such as a cross-encoder reranker Nogueira and Cho (2020)) simply return the highest individually relevant passages, without regard to whether the information in the passages is redundant. Given the premium value on LLM context-window real estate, it’s important to make best use of this limited resource by minimizing redundancy, while maintaining relevance.
@@
A family of methods from the Information Retrieval literature attempts to address the general issue of diversity in retrieved results by introducing a measure that explicitly balances diversity and relevance Carbonell and Goldstein (1998). In this paper, we propose a more principled method, Dartboard, that instead seeks to directly accomplish what previous methods are indirectly aiming for - maximize the total amount of information relevant for a given query in a set of $k$ results. The intuition behind Dartboard is simple - we assume that one passage is the “correct” one for a given query. Our system is allowed $k$ “guesses” and it aims to maximize the relevance score of its most relevant guess. Since the best guess is not known ahead of time, this score is weighted by the probability of that guess being the most relevant. This objective is sufficient to encourage diversity in the guesses. This is because a redundant guess does little to increase the relevance of the most relevant guess.
@@
  • We demonstrate the effectiveness of Dartboard on Retrieval-Augmented Generation Benchmark (RGB) Chen et al. (2023), a closed-domain question answering task. This benchmark consists of a retrieval component, and an end-to-end question-answering component. We show that the Dartboard algorithm, when used as the retrieval component, outperforms all existing baselines at both the component level and at end-to-end level (§3.1).
@@ 2 Dartboard
where $D$ is a distance function. For $d$ dimensional vectors, $A\subseteq\mathbb{R}^{d}$; under some assumptions, we can use a Gaussian kernel for the distance functions. For example, we can set $P\left(T=t|q,\sigma\right)=\mathcal{N}\left(q,t,\sigma\right)$. Thus, our equation becomes:
@@ 2.1 The Dartboard Algorithm
The Dartboard Algorithm aims to maximize Equation 2 given a distance metric. In practice, we can greedily build our set $G$, which works well as it saves us combinatorial search, and allows reuse of previous answers (since the top-$k$ results are a subset of the top-$k+1$ results). We begin by ranking top-$k$ passages $A^{\prime}$ from our initial dataset of passages $A$ using $K$-nearest-neighbors based on cosine similarity. We use a linear search, but sub-linear methods such as FAISS Douze et al. (2024) could also be used for this initial ranking. Our search is a simple greedy optimization method with two changes - (a) we stay in log space to avoid numerical underflow, and (b) we reuse the results ($maxes$) from previous loops to avoid recomputing the maximums. The detailed algorithm is given in Algorithm 1 in Appendix A.1. In Appendix A.3, we also show how to adapt Dartboard to use a cross-encoder based reranker (resulting in two methods called Dartboard crosscoder and Dartboard hybrid), and Appendix A.4 shows that Dartboard generalizes KNN and MMR retrieval algorithms Onal et al. (2015).
@@ 3 Experiments
We tested Dartboard on benchmark datasets from Chen et al. (2023), from which we used two types of closed-domain question answering. In the simple question answering case, a query is answerable from a single passage retrieved from the corpus. For example, consider the query When is the premiere of ‘Carole King & James Taylor: Just Call Out My Name’?. On the other hand, in the information integration case, a query would require multiple passages to be retrieved to answer the query. For example, consider the query Who is the director of ‘Carole King & James Taylor: Just Call Out My Name’ and when is its premiere?. We modified this benchmark for our setup in the following way. The original benchmark contains “positive” and “negative” labeled passages for each query. The positive passages are useful for answering, while the negative ones are related but ineffective in answering the query. Since we are interested in the retrieval component of this task, we merged the positive and negative passages for all queries into a single collection of $11,641$ passages for the $300$ simple question answering test cases and $5,701$ passages for the $100$ information integration test cases. The evaluation is otherwise identical apart from the retrieval component. Note that the innovation of Dartboard is solely on the retrieval component. Therefore, we keep the rest of the RAG pipeline fixed. In particular, we do not modify the prompting of LLMs or try to optimize passage embeddings.
Given a query and the full set of thousands of passage embeddings, we measured both a direct retrieval score and the overall end-to-end performance of the system with the only change being the retrieval algorithm. For the direct retrieval score, we computed the Normalized Discounted Cumulative Gain (NDCG) score Wang et al. (2013) on retrieving any one of the “positive” passages relevant to a specific query. In the information integration case, the positive passages were split into positive ones for each component of the question. Therefore, in this case, we calculated the NDCG score for retrieving at least one positive passage for each component of the query. For the end-to-end score, given an LLM’s response to the query (generated from retrieved passages), we use the same evaluation as Chen et al. (2023), which does a string match of the response on a set of correct answers, marking each response as either correct or incorrect.
@@ 4 Related Work
MMR retrieves documents Carbonell and Goldstein (1998) that are both relevant to the query and dissimilar to previously retrieved documents. It combines a relevance score (e.g., from BM25) with a novelty score that penalizes documents similar to those already retrieved. It have been used extensively for building recommendation systems Xia et al. (2015); Wu et al. (2023) as well as for summarization tasks Agarwal et al. (2022); Adams et al. (2022). However, MMR suffers from few limitations. First is that MMR requires the diversity parameter to control the balance between relevance and novelty. This parameter is often dataset-specific and requires careful tuning, making it impractical for real-world applications. Second is that MMR can favor exact duplicates of previously retrieved documents as they retain a high relevance score while minimally impacting the average novelty score (See Appendix A.7).
KNN retrieves documents based on their similarity to a query embedding Dharani and Aroquiaraj (2013); Bijalwan et al. (2014). While efficient, KNN often suffers from redundancy as nearby documents in the embedding space tend to be semantically similar Taunk et al. (2019). This can lead to a retrieved set dominated by passages conveying the same information with slight variations.
Several recent works have explored incorporating diversity objectives into retrieval models Angel and Koudas (2011); Li et al. (2015); Fromm et al. (2021). These approaches often involve complex optimization functions or require additional training data for diversity estimation. For example, Learning-to-Rank with Diversity methods leverage learning-to-rank frameworks that incorporate diversity objectives directly into the ranking function. This allows for the optimization of both relevance and diversity during the ranking process. However, these approaches often require large amounts of labeled training data for diversity, which can be expensive and time-consuming to obtain Wasilewski and Hurley (2016); Yan et al. (2021). Bandit-based approaches model document selection as a multi-armed bandit problem Hofmann et al. (2011); Wang et al. (2021). The model explores different retrieval strategies and receives feedback based on the relevance and diversity of the retrieved passages. These approaches can be effective but can be computationally expensive for large-scale retrieval tasks.
RAG models have also been extended to incorporate diversity objectives. For example, RAG with Dense Passage Retrieval retrieves a large number of candidate passages Cuconasu et al. (2024); Reichman and Heck (2024); Siriwardhana et al. (2023). It then employs a two-stage selection process: first selecting a diverse subset based on novelty scores, then selecting the most relevant passages from this subset. While effective, this approach requires careful tuning of the selection thresholds.
@@ 6 Limitations
Our experimental results are limited to a single benchmark and a single LLM i.e. ChatGLM Hou et al. (2024). It remains to be seen whether our results would generalize to other benchmarks and LLMs. We plan to investigate this in future work.
@@ A.3 Modification for cross-encoder based reranker
Cross-encoder-based reranking has been shown to outperform embedding-based approaches such as cosine similarity Nogueira and Cho (2020), as it uses the full computational power of a transformer model, rather than being limited to simple vector operations. We have proposed two variations of Dartboard, namely Dartboard Crosscoder and Dartboard Hybrid, based on how we compute the cross-encoder scores for the Gaussian kernels in Equation 2 given in the main paper. For the Dartboard Crosscoder variation, we use the cross-encoder score $C\left(q,t\right)$ before computing the Gaussian kernel for both $\mathcal{N}\left(q,t,\sigma\right)$ and $\mathcal{N}\left(t,g,\sigma\right)$ in Equation 2. Note that the cross-encoder score is asymmetric, so we simply average the two possible ways to compute the cross-encoder score for $\mathcal{N}\left(t,g,\sigma\right)$, i.e., $\frac{1}{2}\left(C\left(t,g\right)+C\left(g,t\right)\right)$. For $\mathcal{N}\left(q,t,\sigma\right)$, we are only interested in the likelihood of $t$ given $q$, so we only use the cross-encoder score $C\left(q,t\right)$.

### v2

# Dartboard: Better RAG using Relevant Information Gain

@@ 1 Introduction
A limitation of transformer-based Large Language Models (LLMs) is that the number of tokens is bounded by the transformer’s context window, which is typically in the thousands. This is often insufficient for representing large texts, such as novels and corporate documentation. A common way to mitigate this constraint is via retrieval augmented generation (RAG), in which a relatively small subset of relevant passages are retrieved from a larger database and inserted into an LLM’s context window (Gao et al. 2024). Typically, this process involves applying a similarity metric, such as cosine similarity, to (precomputed) embeddings of passages and the embedding of a query. Using this metric, many systems then use K-nearest-neighbors or a fast approximation with a vector database such as FAISS (Douze et al. 2024). Importantly, K-nearest-neighbors (Bijalwan et al. 2014) and related methods (such as a cross-encoder reranker (Nogueira and Cho 2020)) simply return the highest individually relevant passages, without regard to whether the information in the passages is redundant. Given the premium value on LLM context-window real estate, it’s important to make best use of this limited resource by minimizing redundancy, while maintaining relevance.
@@
A family of methods from the Information Retrieval literature attempts to address the general issue of diversity in retrieved results by introducing a measure that explicitly balances diversity and relevance (Carbonell and Goldstein 1998). In this paper, we propose a more principled method, Dartboard, that instead seeks to directly accomplish what previous methods are indirectly aiming for - maximize the total amount of information relevant for a given query in a set of $k$ results. The intuition behind Dartboard is simple - we assume that one passage is the “correct” one for a given query. Our system is allowed $k$ “guesses” and it aims to maximize the relevance score of its most relevant guess. Since the best guess is not known ahead of time, this score is weighted by the probability of that guess being the most relevant. This objective is sufficient to encourage diversity in the guesses. This is because a redundant guess does little to increase the relevance of the most relevant guess.
@@
  • We demonstrate the effectiveness of Dartboard on Retrieval-Augmented Generation Benchmark (RGB) (Chen et al. 2023), a closed-domain question answering task. This benchmark consists of a retrieval component, and an end-to-end question-answering component. We show that the Dartboard algorithm, when used as the retrieval component, outperforms all existing baselines at both the component level and at end-to-end level (§3.1).
@@ 2 Dartboard
where $D$ is a distance function. For $d$ dimensional vectors, $A\subseteq\mathbb{R}^{d}$; under some assumptions, we can use a Gaussian kernel for the distance functions. For example, we can set $P\left(T=t|q,\sigma\right)=\mathcal{N}\left(q,t,\sigma\right)$, and we can set $D\left(t|g\right)\propto-\mathcal{N}\left(t,g,\sigma\right)$. Thus, our equation becomes:
@@ 2.1 The Dartboard Algorithm
The Dartboard Algorithm aims to maximize Equation 2 given a distance metric. In practice, we can greedily build our set $G$, which works well as it saves us combinatorial search, and allows reuse of previous answers (since the top-$k$ results are a subset of the top-$k+1$ results). We begin by ranking top-$k$ passages $A^{\prime}$ from our initial dataset of passages $A$ using $K$-nearest-neighbors based on cosine similarity. We use a linear search, but sub-linear methods such as FAISS (Douze et al. 2024) could also be used for this initial ranking. Our search is a simple greedy optimization method with two changes - (a) we stay in log space to avoid numerical underflow, and (b) we reuse the results ($maxes$) from previous loops to avoid recomputing the maximums. The detailed algorithm is given in Algorithm 1 in Appendix A.1. In Appendix A.3, we also show how to adapt Dartboard to use a cross-encoder based reranker (resulting in two methods called Dartboard crosscoder and Dartboard hybrid), and Appendix A.4 shows that Dartboard generalizes KNN and MMR retrieval algorithms (Onal, Altingovde, and Karagoz 2015).
@@ 3 Experiments
We tested Dartboard on benchmark datasets from (Chen et al. 2023), from which we used two types of closed-domain question answering. In the simple question answering case, a query is answerable from a single passage retrieved from the corpus. For example, consider the query When is the premiere of ‘Carole King & James Taylor: Just Call Out My Name’?. On the other hand, in the information integration case, a query would require multiple passages to be retrieved to answer the query. For example, consider the query Who is the director of ‘Carole King & James Taylor: Just Call Out My Name’ and when is its premiere?. We modified this benchmark for our setup in the following way. The original benchmark contains “positive” and “negative” labeled passages for each query. The positive passages are useful for answering, while the negative ones are related but ineffective in answering the query. Since we are interested in the retrieval component of this task, we merged the positive and negative passages for all queries into a single collection of $11,641$ passages for the $300$ simple question answering test cases and $5,701$ passages for the $100$ information integration test cases. The evaluation is otherwise identical apart from the retrieval component. Note that the innovation of Dartboard is solely on the retrieval component. Therefore, we keep the rest of the RAG pipeline fixed. In particular, we do not modify the prompting of LLMs or try to optimize passage embeddings.
Given a query and the full set of thousands of passage embeddings, we measured both a direct retrieval score and the overall end-to-end performance of the system with the only change being the retrieval algorithm. For the direct retrieval score, we computed the Normalized Discounted Cumulative Gain (NDCG) score (Wang et al. 2013) on retrieving any one of the “positive” passages relevant to a specific query. In the information integration case, the positive passages were split into positive ones for each component of the question. Therefore, in this case, we calculated the NDCG score for retrieving at least one positive passage for each component of the query. For the end-to-end score, given an LLM’s response to the query (generated from retrieved passages), we use the same evaluation as (Chen et al. 2023), which does a string match of the response on a set of correct answers, marking each response as either correct or incorrect.
@@ 4 Related Work
MMR retrieves documents (Carbonell and Goldstein 1998) that are both relevant to the query and dissimilar to previously retrieved documents. It combines a relevance score (e.g., from BM25) with a novelty score that penalizes documents similar to those already retrieved. It have been used extensively for building recommendation systems (Xia et al. 2015; Wu, Wang, and Ma 2023) as well as for summarization tasks (Agarwal, Xu, and Grabmair 2022; Adams, Suri, and Chali 2022). However, MMR suffers from few limitations. First is that MMR requires the diversity parameter to control the balance between relevance and novelty. This parameter is often dataset-specific and requires careful tuning, making it impractical for real-world applications. Second is that MMR can favor exact duplicates of previously retrieved documents as they retain a high relevance score while minimally impacting the average novelty score (See Appendix A.7).
KNN retrieves documents based on their similarity to a query embedding (Dharani and Aroquiaraj 2013; Bijalwan et al. 2014). While efficient, KNN often suffers from redundancy as nearby documents in the embedding space tend to be semantically similar (Taunk et al. 2019). This can lead to a retrieved set dominated by passages conveying the same information with slight variations.
Several recent works have explored incorporating diversity objectives into retrieval models (Angel and Koudas 2011; Li et al. 2015; Fromm et al. 2021). These approaches often involve complex optimization functions or require additional training data for diversity estimation. For example, Learning-to-Rank with Diversity methods leverage learning-to-rank frameworks that incorporate diversity objectives directly into the ranking function. This allows for the optimization of both relevance and diversity during the ranking process. However, these approaches often require large amounts of labeled training data for diversity, which can be expensive and time-consuming to obtain (Wasilewski and Hurley 2016; Yan et al. 2021). Bandit-based approaches model document selection as a multi-armed bandit problem (Hofmann et al. 2011; Wang, Jia, and Wang 2021). The model explores different retrieval strategies and receives feedback based on the relevance and diversity of the retrieved passages. These approaches can be effective but can be computationally expensive for large-scale retrieval tasks.
RAG models have also been extended to incorporate diversity objectives. For example, RAG with Dense Passage Retrieval retrieves a large number of candidate passages (Cuconasu et al. 2024; Reichman and Heck 2024; Siriwardhana et al. 2023). It then employs a two-stage selection process: first selecting a diverse subset based on novelty scores, then selecting the most relevant passages from this subset. While effective, this approach requires careful tuning of the selection thresholds.
@@ 6 Limitations
Our experimental results are limited to a single benchmark and a single LLM i.e. ChatGLM (Hou et al. 2024). It remains to be seen whether our results would generalize to other benchmarks and LLMs. We plan to investigate this in future work.
@@ A.3 Modification for cross-encoder based reranker
Cross-encoder-based reranking has been shown to outperform embedding-based approaches such as cosine similarity (Nogueira and Cho 2020), as it uses the full computational power of a transformer model, rather than being limited to simple vector operations. We have proposed two variations of Dartboard, namely Dartboard Crosscoder and Dartboard Hybrid, based on how we compute the cross-encoder scores for the Gaussian kernels in Equation 2 given in the main paper. For the Dartboard Crosscoder variation, we use the cross-encoder score $C\left(q,t\right)$ before computing the Gaussian kernel for both $\mathcal{N}\left(q,t,\sigma\right)$ and $\mathcal{N}\left(t,g,\sigma\right)$ in Equation 2. Note that the cross-encoder score is asymmetric, so we simply average the two possible ways to compute the cross-encoder score for $\mathcal{N}\left(t,g,\sigma\right)$, i.e., $\frac{1}{2}\left(C\left(t,g\right)+C\left(g,t\right)\right)$. For $\mathcal{N}\left(q,t,\sigma\right)$, we are only interested in the likelihood of $t$ given $q$, so we only use the cross-encoder score $C\left(q,t\right)$.